OpenStudy (laurennarena):

Dan bought x pounds of potatoes for $0.85 per pound and y pounds of grapes for $1.29 per pound. The total cost was less than $5. Which inequality represents this purchase? 1.29x + 0.85y < 5 1.29x + 0.85y > 5 0.85x + 1.29y > 5 0.85x + 1.29y < 5

OpenStudy (doc.brown):

What's your first thought?

OpenStudy (doc.brown):

\(x\) pounds of potatoes at a rate of $0.85 per pound of potatoes\[x\text{ lb}\times\frac{$0.85}{\text{lb}}\]

OpenStudy (doc.brown):

\[x\text{ lb}\times\frac{$0.85}{\text{lb}}\]is the same as\[\frac{x\text{ lb}\times$0.85}{\text{lb}}\]which is the same as\[x\times$0.85\times\frac{\text{lb}}{\text{lb}}\]or\[$0.85x\times\frac{\text{lb}}{\text{lb}}\]